OpenStudy (stormoshawty):

Find the slope for 5x + 3y = -2

OpenStudy (stormoshawty):

@shamil98

OpenStudy (anonymous):



OpenStudy (shamil98):

5x + 3y = -2 You want to get this in slope-intercept form , y=mx +b so solve for y. 3y = -5x - 2 divide by 3 y = -5/3 x - 2/3 slope = m = -5/3

OpenStudy (stormoshawty):

I love you guys. Thank you.

OpenStudy (anonymous):

you can also try to remember that the slope of \[\large ax+by=c\] is \[\large m=-\frac{a}{b}\] may save a step or two

OpenStudy (anonymous):

thank you @satellite73 i didn't know that :D